### v1

# Numerical simulations of density perturbation and gravitational wave production from cosmological first-order phase transition

Abstract
We conducted three-dimensional lattice simulations to study the density perturbation and generation of gravitational waves (GWs) through first-order phase transition (FOPTs). Our findings reveal that for the case of $\beta/H=6$, density perturbations successfully collapse into PBHs. Moreover, smaller values of $\beta/H$ are more conducive to PBH formation, while the strength of the phase transition $\alpha$ has a relatively minor impact on PBH production. We find that for $\alpha>1$, the dominant contribution to the density perturbation comes from the delay of vacuum decay, whereas for $\alpha<1$, the bubble walls’ moving forward behavior becomes the primary source. Additionally, the power spectrum of density perturbations generated by the phase transition exhibits a slope of $k^{3}$ at small wavenumbers and $k^{-1.5}$ at large wavenumbers. Furthermore, we calculated the GW power spectra, and the simulation results show a slope of $k^{3}$ at small wavenumbers and $k^{-2}$ at large wavenumbers. Our numerical simulations confirm that slow PTs can produce PBHs and provide predictions for the GW power spectrum, offering crucial theoretical support for GW detection.
first order phase transition, density perturbation
I INTRODUCTION
First-order phase transition (FOPT) is one important source of stochastic gravitational wave (GW) background Hindmarsh et al. (2021). During the FOPT, the anisotropic stresses produced by the bubble collisions Cutting et al. (2018, 2021), sound waves Hindmarsh et al. (2015, 2017, 2014); Cutting et al. (2020, 2022); Giblin and Mertens (2014), and magnetohydrodynamic (MHD) turbulence Caprini et al. (2009); Kahniashvili et al. (2008, 2010); Caprini et al. (2016); Wansleben et al. (2019) generate GW radiations. The resulting GW spectrum encodes information about the properties of the phase transition, such as its strength, duration, and characteristic energy scale. Observing these GWs with current and future detectors provides a direct probe of the underlying particle physics responsible for the FOPT Athron et al. (2024); Caldwell et al. (2022); Bian et al. (2021), including extensions to the Standard Model, such as additional scalar fields or symmetry-breaking mechanisms.
In the early universe, FOPT occurs when the cosmic temperature drops below the critical temperature, triggering vacuum bubble nucleation in the false vacuum background. The subsequent expansion, collision, and percolation of these bubbles release significant latent heat and can generate large density fluctuations from random generation of vacuum bubbles Liu et al. (2022). If these fluctuations exceed a critical threshold, they may collapse under their own gravity to form Primordial black holes (PBHs) Hawking et al. (1982); Crawford and Schramm (1982); Kodama et al. (1982); Johnson et al. (2012); Khlopov et al. (1998). PBH is hypothesized to form in the early universe due to large density perturbations, offer a unique window into the physics of the early cosmos. Unlike stellar-mass black holes formed from collapsing stars, PBHs can span a wide range of masses, from sub-planetary to stellar and even supermassive scales, depending on their formation mechanism and the underlying physics. They have attracted considerable attention as potential candidates for dark matter, sources of high-energy phenomena, and contributors to the gravitational wave background observed by detectors such as LIGO-Virgo Sasaki et al. (2016); Bird et al. (2016); Clesse and García-Bellido (2017); Hütsi et al. (2021); Hall et al. (2020); Franciolini et al. (2022); He et al. (2024), and the planned LISA Armano et al. (2016); Amaro-Seoane et al. (2017); Seoane et al. (2023), Taiji Hu and Wu (2017); Ruan et al. (2020); Wu et al. (2021) as well as TianQin mission Luo et al. (2016, 2020); Mei et al. (2021). There are currently many proposed mechanisms for the formation of PBHs, such as inflaton ultraslow-roll Carr and Lidsey (1993); Ivanov et al. (1994), the collapse of cosmic strings Hawking (1989); Caldwell and Casper (1996); Jenkins and Sakellariadou (2020); Blanco-Pillado et al. (2021), domain walls Vachaspati (2017); Ferrer et al. (2019); Gelmini et al. (2023a, b), scalar condensates Dolgov and Silk (1993); Kasai et al. (2022); Cotner and Kusenko (2017); Martin et al. (2020), or processes in a dissipative dark sector Chang et al. (2019); Flores and Kusenko (2021); Domènech et al. (2023); Chakraborty et al. (2022).
The mass spectrum and abundance of PBHs produced through FOPT depend sensitively on the dynamics of the phase transition, such as the nucleation rate, bubble wall velocity, and phase transition strength of the transition Lewicki et al. (2024). In this work, we perform numerical simulations to investigate the density perturbation in the early universe, originating from delayed vacuum tunneling. In this scenario, specific regions of the universe undergo delayed decay of the false vacuum. As the radiation energy density decreases over time while the vacuum energy remains constant, this leads to the formation of localized energy overdensities. If these overdensities exceed a critical threshold, they may collapse into PBHs. Additionally, we compare the effects of different phase transition rates and strengths on the simulation results and further simulate the GW spectra generated by bubble expansion and collisions during the FOPT. These findings provide valuable insights into the dynamics of the early universe. For convenience, we adopt the reduced Planck mass $m_{pl}=1$ and Gravitational constant $G=1$ throughout this study.
II Density perturbations FROM FOPTs
We consider a phase transition involving the scalar field $\phi$, and the associated potential can be expressed aspotential Cutting et al. (2018, 2021)
$V(\phi)=\frac{1}{2}M^{2}\phi^{2}+\frac{1}{3}\delta\phi^{3}+\frac{1}{4}\lambda\phi^{4}+\Delta V,$ (1)
where $\Delta V$ is a constant used to adjust the potential’s minimum value to zero. During the transition, vacuum bubbles nucleate with an exponential rate expressed as Coleman (1977); Enqvist et al. (1992)
$\Gamma(t)=\Gamma_{0}e^{\beta t},$ (2)
where $\Gamma_{0}$ is the initial nucleation rate, and $\beta$ is a key parameter that dictates the rate’s increase over time. After nucleation, the field evolves according to the Klein-Gordon equation in an expanding universe Child and Giblin (2012)
$\ddot{\phi}+3H\dot{\phi}-\frac{\nabla^{2}\phi}{a^{2}}+\frac{\partial V(\phi)}{\partial\phi}=0,$ (3)
the spatially averaged energy density and pressure can be split into four componentsare Adshead et al. (2020)
$\rho_{tot}$ $=\rho_{K}+\rho_{G}+\rho_{V}+\rho_{r}$ (4)
$P$ $=\rho_{K}-\frac{1}{3}\rho_{G}-\rho_{V}+\frac{1}{3}\rho_{r},$ (5)
where $\rho_{K}=\frac{\phi^{\prime}}{2a^{2}}$, $\rho_{G}=\frac{(\partial_{i}\phi)^{2}}{2a^{2}}$, $\rho_{V}$, and $\rho_{r}$ are the spatially averaged kinetic energy density, gradient energy density, potential energy density, and energy density of background radiation, respectively. The PT strength is defined as $\alpha=\Delta V/\rho_{r}$, which represents the ratio of vacuum energy released to the background radiation energy density. The continuity equation can be written as
$\frac{{\mathrm{d}(\rho_{K}+\rho_{G}+\rho_{r})}}{\mathrm{d}t}+3\mathcal{H}(2\rho_{K}+\frac{2}{3}\rho_{G}+\frac{4}{3}\rho_{r})=-\frac{{\mathrm{d}\rho_{V}}}{\mathrm{d}t}\;.$ (6)
Where, $\mathcal{H}=a^{\prime}/a$ is the conformal Hubble parameter related to the usual Hubble parameter via $\mathcal{H}=aH$ with $a^{\prime}\equiv\partial a/\partial t$. The corresponding Friedmann’s equation is given as
$\mathcal{H}^{2}=\frac{8\pi}{3}a^{2}\rho_{tot}.$ (7)
During the PT, asynchronous vacuum tunneling results in regions within each Hubble-sized volume where the false vacuum decay is delayed. In an expanding universe, while the radiation energy density decreases, the vacuum energy remains constant. This delay in false vacuum decay increases the local energy density in these regions. During the mid-to-late stages of a FOPT, density perturbations on the Hubble scale cross the horizon, leading to significant fluctuations with
$\delta=\frac{\rho_{\rm{inside}}-\rho_{\rm{outside}}}{\rho_{\rm{outside}}}\;,$ (8)
where $\rho_{inside}$ and $\rho_{outside}$ represent the energy densities inside and outside the delayed decay region, respectively.
The power spectrum of these perturbations is crucial for understanding the characteristics of density fluctuations. The dimensionless power spectrum can be obtained by directly differentiating the variance at the Hubble scale
$\mathcal{P_{\delta}}(k,t)=\frac{\partial\rho_{\delta}}{\partial lnk},$ (9)
where $\rho_{\delta}=(\frac{\rho-\bar{\rho}}{\bar{\rho}})^{2}$ , $\delta_{H}$ is the standard deviation of $\rho_{\delta}=(\frac{\rho-\bar{\rho}}{\bar{\rho}})^{2}$ at the Hubble horizon scale. It can be expressed as
$\delta_{H}^{2}(k,t)=\int\frac{dk}{k}\mathcal{P_{\delta}}(k,t)\;.$ (10)
The curvature perturbation power specturm can be obtained from the following equation Josan et al. (2009); Ando et al. (2018)
$\mathcal{P}_{\delta}(k,t)=\frac{16}{3}\left(\frac{k}{aH}\right)^{2}j_{1}^{2}(k/\sqrt{3}aH)\mathcal{P}_{\mathcal{R}}(k,t),$ (11)
where $j_{1}$ is a spherical Bessel function.
III GW power specturm definition
During a FOPT, vacuum bubbles are generated throughout space via quantum tunneling and then rapidly expand. As more and more bubbles form and collide, the initially uniform spatial structure is disrupted, resulting in an asymmetric energy distribution. The walls of these bubbles carry high energy, and when they collide, this energy is intensely released, creating asymmetric disturbances. This asymmetry breaks the spherical symmetry of the system, causing violent oscillations in spacetime curvature. These oscillations propagate as waves throughout the universe, producing GWs.
GWs only require considering the transverse and traceless components. Therefore, the energy-momentum tensor for GWs is Hindmarsh et al. (2015)
$T_{\mu\nu}^{\text{GW}}=\frac{1}{32\pi G}\langle\partial_{\mu}h_{ij}\partial_{\nu}h_{ij}\rangle,$ (12)
where $h_{ij}$ represents the small perturbation of the metric. Therefore, the GW power spectrum is then
$\mathcal{P}_{gw}(t)=\frac{1}{24\pi^{2}\mathcal{H}^{2}V}\sum_{i,j}\int\mathrm{d}\Omega\left|\mathbf{k}\right|^{3}\left|h_{ij}^{\prime}(k)\right|^{2}$ (13)
where $V$ is the simulation volume. What is important in our simulation is that the momentum $k$ has a minimum infrared cut-off because of the discrete lattice.
IV NUMERICAL RESULTS
In this simulation, we consider 64 Hubble volumes and define the last 16 to form a bubble center as postponed Hubble volumes. We calculate the energy fluctuation of each Hubble volume whose decay of false vacuum is postponed. In a rapid phase transition process, every Hubble volume can contain at least one bubble center in the end. Hence, we artificially choose that the Hubble volume getting a bubble after the time that $1/4$(maybe $1/e$ is better) of all Hubble volumes have not contained any bubble center is called the postponed Hubble volume. In a slow phase transition, due to the slow bubble nucleation rate. In this case, we can simply define the postponed Hubble volume is the region that do not contain any bubble center. Typically, in this simulation, we employ a lattice of $N^{3}=512^{3}dx$ to model a region with $LH=4$, and $H$ being the Hubble parameter. Corresponding to a total of $4^{3}$ Hubble volumes. The initial bubble radius and wall thickness are set to $12dx$ and $4dx$, respectively. The mean bubble separation is obtained as $R^{*}=(L^{3}/N_{b})^{\frac{1}{3}}$ with $N_{b}$ being the number of the generated bubbles during the FOPT processes. For the vacuum bubble nucleation probability given by Eq. (2), we explore multiple values of $\alpha=0.5,1,5,10$ and $\beta/H=6,8,10,12$. The spatial and temporal lattice spacings are defined as $\mathrm{d}x=L/N$ and $\mathrm{d}t=\mathrm{d}x/5$, respectively. The numerical calculations is are conducted with a code developed based on pystella Weiner (2021).
Figure 1: EOS evolution for weak phase transition strength (left, solid line $\alpha$=0.5, dashed line $\alpha$=1) and strong phase transition strength (right, solid line $\alpha$=5, dashed line $\alpha$=10). Four colors represent different $\beta/H$ values. In all cases, $\omega$ transitions from $\omega\to-1$(vacuum energy domination) to a radiation-dominated state with $\omega=1/3$. A smaller $\alpha$ results in a longer transition time to reach $\omega=1/3$.
The evolution of the average energy density of each component follows the evolution described by the continuity equation Eq. (6) and the Friedmann equation Eq. (7). As the phase transition progresses, the vacuum bubbles expand, converting the vacuum energy $\rho_{V}$ into the bubble wall energy $\rho_{w}$ (including the kinetic and gradient energy of the wall) and the background radiation energy $\rho_{r}$. Initially, bubble wall energy increases, reaching a peak before decreasing, which corresponds to changes in the phase transition rate. When the PT rate reaches its maximum, numerous vacuum bubbles form in space, with the earliest bubbles fully expanded. Consequently, the wall energy reaches its peak. As bubbles collide and merge, the bubble wall energy gradually converts into background radiation energy, eventually approaching zero. Since $\rho_{r}\propto a(t)^{-4}$, the background radiation energy density also tends towards zero after the phase transition is complete. Subsequently, the Equation of State (EoS), which is computed by $\omega=P/\rho_{tot}$, is shown in Fig. 1. It can be seen that in the beginning, the universe was dominated by vacuum energy, also known as dark energy, with a value of $\omega\to-1$. As the vacuum energy is continuously converted into the radiation energy, $\omega$ continues to increase. After the phase transition of the entire space is completed, the universe enters a radiation-dominated stage, and ultimately $\omega=1/3$.
Figure 2: Evolution diagram of $\delta_{H}(t)=\int\frac{dk}{k}\mathcal{P_{\delta}}(t)$ during FOPT process. The panels are arranged in a $2\times 2$ grid: $\alpha=0.5$ (top-left), $\alpha=1$ (top-right), $\alpha=5$ (bottom-left), and $\alpha=10$ (bottom-right). The contribution from smaller scales ($kR^{*}>1.3$) has been droped.
For the density perturbation, we compute the value of $\delta_{H}$ using Eq. (10), and the corresponding evolution is presented in Fig. 2. Where, the $\delta_{H}$ for PTs of $\alpha=0.5,1$ tends to reach the maximum when $t/R^{*}\sim 1$ mostly come from moving forward of the bubble walls. Differently, the $\delta_{H}$ reach maximum at around $t/R^{*}\sim 0.5$ for strong PTs scenarios with $\alpha=5,10$ sourced from the delay of vacuum decay in different Hubble volumes, and the second peak comes from the energy transformation in the simulation volume due to the vacuum bubble walls’ moving forward behavior. Therefore, we present $\delta_{H}$ as a function of $\alpha$ for different $\beta/H$ at the two peak moments $t/R^{*}=0.5,1.2$ and plot it in Fig. 3. The left plot demonstrates that the $\delta_{H}$ increases as the PT strength increases. While, the right plot show that generally the $\delta_{H}$ increase before $\alpha\leq 1$ and decrease after $\alpha\geq 1$.
Figure 3: For $\beta/H=6,8,10,12$, the left panel shows $\delta_{H}$ at the first peak $t/R^{*}=0.5$, while the right panel presents $\delta_{H}$ at the second peak $t/R^{*}=1.2$.
Figure 4: The power spectra of density perturbations for different $\beta/H$ values during the mid-stage of the phase transition. The $\alpha$ from top-left to bottom-right are 0.5, 1, 5, and 10, respectively.
Fig. 4 shows the power spectrum of density perturbations at the completion of the phase transition. For $log_{10}[kR^{*}]\textless 0.5$, the slope follows $k^{3}$, while for $0.8\textless log_{10}[kR^{*}]\textless 1.3$, the slope is $k^{-1.5}$. Notably, no clear correlation is observed between the power spectrum and the parameters $\beta/H$ or $\alpha$, suggesting that the density perturbation spectrum primarily depends on the intrinsic dynamics of the phase transition rather than its specific strength or nucleation rate. We use a function $f_{\delta}$ to fit $P_{\delta}=f_{\delta}k^{3,-1.5}$ for infra-red (IR) and ultra-violet (UV) wavenumber regions
$f_{\delta}(\alpha,\beta/H)=A(\beta/H)^{n}\alpha^{m}\;.$ (14)
For the cases where $\alpha<1$, the fitting results are $A=2.24$, $n=-0.81$, and $m=0.35$. And, for the eight cases where $\alpha>1$, the fitting results are $A=0.14$, $n=0.62$, and $m=-0.39$.
Figure 5: The energy density spectra of GWs for different $\beta/H$ values at the late stage of the phase transition. The $\alpha$ from top-left to bottom-right are 0.5, 1, 5, and 10, respectively.
When vacuum bubbles expand and collide, the large amount of energy carried by their walls is released violently and creates asymmetric disturbances, making it a source of GWs Falkowski and No (2013); Konstandin and Servant (2011). In Fig. 5, we show the energy density spectrum of GW for FOPT. For $log_{10}[kR^{*}]\textless 0.5$, the GW power spectrum exhibits a steep rise with a slope of $k^{3}$, which is a little bit different for the $\alpha=0.5$ case and the discrepancy may be caused by the numerical fault. In the range of $0.8\textless log_{10}[kR^{*}]\textless 1.3$, the slope is $k^{-2}$, Moreover, our analysis reveals that the overall shape and amplitude of the power spectrum show little sensitivity to the parameters $\alpha$ and $\beta/H$. Meanwhile, according to our calculations for the $\alpha$ is $0.5$ and $1$, the magnitude is larger when $\beta/H$ is smaller, which would disppear at $\alpha=$ $5$ and $10$.
If this overdensity exceeds the threshold $\delta_{c}$, the Hubble patch collapses into a PBH Musco et al. (2005); Harada et al. (2013). During the the mid-to-late phase transition period, the value of threshold is within the range of 0.40 to 0.67 Musco (2019); Escrivà et al. (2020), in this work, we select commonly used $\delta_{c}=0.45$ Jedamzik and Niemeyer (1999); Green et al. (2004); Musco et al. (2005); Musco and Miller (2013); Harada et al. (2015). It is worth mentioning that the condition where the overdensity $\delta$ exceeds the critical value $\delta_{c}$ generally occurs before the vacuum energy fully decays in the overly dense regions. Furthermore, the results indicate that PBH formation occurs for slow phase transitions with $\beta/H=6$, see Appendix B and Appendix. C for more details during the PT processes. Refs. Liu et al. (2022); Kawana et al. (2023); Cai et al. (2024) suggest that a smaller $\beta/H$ leads to a higher PBH abundance, which aligns with our simulation results. Regarding the effect of $\alpha$, these studies conclude that PBH abundance increases as $\alpha$ increases. However, our simulation results indicate that $\alpha$ has little impact on PBH abundance, making it difficult to draw a definitive conclusion. In the future, with improved simulation precision, a more accurate assessment of $\alpha$’s effect may be possible.
V CONCLUSION AND DISCUSSION
In this study, we performed three-dimensional numerical simulations to investigate the density perturbation and GWs production during FOPTs. Our simulations show that regions undergoing delayed vacuum decay can indeed develop local overdensities, and that smaller values of $\beta/H$ facilitate the formation of such overdense regions. In our simulations, variations in $\alpha$ did not significantly affect the formation of high-density regions, suggesting that higher-precision simulations may be needed to clarify this effect. Additionally, in the power spectrum of density perturbations at the completion of the phase transition, the slope follows $k^{3}$ for $\log_{10}[kR^{*}]\textless 0.5$, while in the range $0.8<\log_{10}[kR^{*}]\textless 1.3$, the slope is $k^{-1.5}$. We found that for $\alpha\textless 1$, $\delta_{H}$ is positively correlated with $\alpha$, which is consistent with the results given in Ref. Liu et al. (2023).
We observed two peaks in the density perturbation. The first peak is caused by the delay of vacuum decay in different Hubble volumes, while the second peak originates from the moving forward behavior of the bubble walls. Moreover, for $\alpha>1$, the peak due to the delay of vacuum decay dominates, whereas for $\alpha<1$, the peak arising from the moving forward behavior of the bubble walls becomes dominant. With the fitted curvature perturbation formula, for latter case with the same PT temperature, the CMB spectral distortion considered would yield more strong limit on the PT strength $\alpha$ and weaker limit on the PT inverse duration $\beta/H$ than given by Ref. Liu et al. (2023). The situation is in contrast with the strong PT scenarios of $\alpha>1$. For more detailed on the BBN and $\Delta N_{eff}$ constraints on the PT parameters, we refer to Refs.Bai and Korwar (2022); Deng and Bian (2023). For the superhorizon fluctuations induced by delayed vacuum decay at low scale PTs with $T_{PT}\lesssim$GeV, see Ref. Elor et al. (2024).
Furthermore, the GW spectra were shown to have the slope following $k^{3}$ and $k^{-2}$ at small and large wave numbers. We did not observe a significant dependence on $\alpha$ and $\beta/H$ for these slope. Additionally, when analyzing the GW power spectrum at different stages of the phase transition, we did not find any notable changes in its slope. In the simulations with $\alpha=0.5$ and $\alpha=1$, we found that as $\beta/H$ decreases, the magnitude is larger. This finding is consistent with the conclusions of Ref. Lewicki et al. (2024).
Acknowledgements.
We thank Jing Liu for fruitful discussion. This work is supported by the National Key Research and Development Program of China under Grant No. 2021YFC2203004, and by the National Natural Science Foundation of China (NSFC) under Grants Nos. 12322505, 12347101. We also acknowledges Chongqing Talents: Exceptional Young Talents Project No. cstc2024ycjh-bgzxm0020 and Chongqing Natural Science Foundation under Grant No. CSTB2024NSCQ-JQX0022.
Appendix A Phase transition process
The evolution of the scalar field $\phi$ is depicted in Fig. 6. It can be observed that the mean value of the field gradually increases from zero, corresponding to the nucleation and expansion of vacuum bubbles in space. Initially, the growth rate of the field’s mean value accelerates before slowing down and eventually reaching a plateau at zero. Moreover, as $\alpha$ increases and $\beta/H$ becomes larger, the completion time of the phase transition, characterized by $t/R^{*}$, decreases. This is expected, as a stronger phase transition facilitates a more rapid completion of the process. The time of phase transition completion and the corresponding number of bubbles are presented in Tab. 1.
Figure 6: The mean field evolution under weak phase transition strength (left, solid line $\alpha$=0.5, dashed line $\alpha$=1) and strong phase transition strength (right, solid line $\alpha$=5, dashed line $\alpha$=10). Four colors represent different $\beta/H$.
Table 1: Bubble parameters for different values of $\beta/H$ and $\alpha$. Each cell contains two quantities: the bubble number $N_{b}$ of the generated bubbles during the FOPT processes, the time $t/R^{*}$ required to reach $\phi/\phi_{b}$=0.999.
$\alpha\backslash\beta/H$ 6 8 10 12
$0.5$ $(281,3.57)$ $(391,4.12)$ $(521,4.67)$ $(601,5.02)$
$1$ $(296,2.71)$ $(397,3.15)$ $(498,3.56)$ $(610,3.93)$
$5$ $(289,1.83)$ $(388,2.13)$ $(481,2.42)$ $(601,2.72)$
$10$ $(278,1.70)$ $(387,2.01)$ $(500,2.30)$ $(598,2.54)$
As an illustration in Fig. 7, we present bubble slices with $\beta/H=6$ and $\alpha=5$, which visually demonstrate the process of bubble expansion and phase transition. Clearly, one could have more energy transormation from bubble walls’s moving forward with the proceeding of the PT. For more details on the energy evolutions of different components, see Fig. 8.
Figure 7: The bubble slice diagram for $\beta/H$ = 6 and $\alpha$ = 5. The dashed lines are presented to devide the simulation box to 64 Hubble patches.
Figure 8: The energy density $\rho$ under weak phase transition strength (left, $\alpha=1$, solid line $\beta/H=6$, dashed line $\beta/H=10$) and strong phase transition strength(right, $\alpha=10$ solid line $\beta/H=6$, dashed line $\beta/H=10$).
Appendix B Evolution of the density perturbation
Using the Press-Schechter formalismPress and Schechter (1974); Ando et al. (2018); Yoo et al. (2018), we obtain an alternative expression for $\delta_{H}(t)$
$\delta_{H}(t)=\int\frac{dk}{k}W^{2}(k)\mathcal{P_{\delta}}(t),$ (15)
Here, $W(k)$ is the window function, which in this case is chosen to be Gaussian
$W(k)=e^{-k^{2}/\mathcal{H}^{2}}\;.$ (16)
The results for different phase transition parameters are shown in Fig. 9, which demonstrate a quite similar behavior as in Fig. 2.
Figure 9: Evolution diagram of $\delta_{H}$ during FOPT process, with $\delta_{H}(t)=\int\frac{dk}{k}W^{2}(k)\mathcal{P_{\delta}}(t)$ where the contribution from smaller scales ($kR^{*}>1.3$) has been neglected. The $\alpha$ from left to right are 0.5, 1, 5, 10, respectively.
Figure 10: $\delta$ of the last 16 the last 16 delayed-decayed regions. The $\beta/H$ from left to right are 6, 8, 10, and 12, the $\alpha$ from top to bottom are 0.5, 1, 5, and 10, respectively.
We numerically analyze the evolution of local energy overdensities $\delta$ induced by delayed vacuum decay during FOPTs in the early universe. Specifically, we examine how different phase transition parameters, $\alpha$ and $\beta/H$, affect $\delta$ and whether these overdense regions can collapse into PBHs. As shown in the Fig. 10, due to the phase transition in the background universe, the vacuum energy gradually converts into the radiation energy, resulting in a lower energy density in the delay region compared to the background universe, and $\delta$ gradually decreases. When the delay region begins to undergo a phase transition, $\delta$ gradually increases. After the energy density disturbance reaches its maximum value, the content of the wall gradually decreases, and the background radiation energy is also averaged throughout the entire space, so $\delta$ gradually approaches zero. In addition, it can be observed that $\alpha$ has little effect on energy density perturbations, while its dependence on $\beta/H$ is extremely strong. The smaller the value of $\beta/H$, the larger the energy density perturbation. When $\beta/H$ = 6, the perturbation reaches the critical threshold $\delta_{c}$ = 0.45, ultimately collapsing to form PBHs. Our simulation results show that for larger $\beta/H$ (faster phase transitions), the probability of $\delta$ exceeding the critical threshold $\delta_{c}=0.45$ is significantly reduced. This trend suggests that PBH formation is suppressed in stronger and faster phase transitions. This phenomenon can be understood through the following mechanisms. When $\alpha$ is large, a greater fraction of vacuum energy is released during the phase transition, leading to a more uniform redistribution of energy and rapid smoothing of density fluctuations. As a result, compared to smaller $\alpha$ cases, the likelihood of forming sufficiently large $\delta$ to trigger PBH formation is reduced. Furthermore, when $\beta/H$ is large, the phase transition proceeds rapidly, with vacuum bubbles nucleating and merging within a short timescale. This results in a more homogeneous phase transition across space, reducing the formation of localized energy overdensities.
Appendix C GW power spectra
Figure 11: From left to right, the GW power spectra correspond to $\beta/H$=6, 8, 10, and 12 in sequence. The $\alpha$ from top to bottom are 0.5, 1, 5, and 10, respectively. Each subplot illustrates the energy density spectrum of GWs at the late stage of the phase transition, where the four solid lines in different colors represent the GW power spectra for different values of $\omega$.
We compute the GW (GW) power spectrum generated during FOPTs and investigate its dependence on different values of the equation of state parameter $\omega$. Our results indicate that when $\alpha$ is 5 and 10, $\beta/H$ is 6 and 8, the slope of the GW power spectrum increases as omega increases, which do not change a lot for other circumstances. The GW power spectrum remains consistent in shape, with no significant variations observed. This suggests that the GW spectrum produced by bubble wall collisions during FOPTs is weakly dependent on $\omega$. However, we note that our conclusions regarding the dependence of the GW spectrum on $\omega$ differ from those reported in Refs. Domènech et al. (2020); Cai et al. (2020), requiring further investigation to clarify this issue.